OpenStudy (accidentalaichan):

Which of the following ordered pairs is a solution to the inequality?

OpenStudy (accidentalaichan):

\[y \ge 4x - 2\]

OpenStudy (accidentalaichan):

A. (0, −4) B. (0, 0) C. (0, −3) D. (0, −5)

OpenStudy (e.mccormick):

Well, do you know how to find the area that is the solution to an inequality?

OpenStudy (e.mccormick):

You can do it that way, or you can just try them out and see which works.

OpenStudy (accidentalaichan):

okay, so I did my inverse operations, adding 2 to each side \[2y \ge 4x\] But now I'm lost

OpenStudy (e.mccormick):

How did you get 2y?

OpenStudy (accidentalaichan):

I added 2 to each side. It wasn't right, was it?

OpenStudy (e.mccormick):

That would be 2+y, not 2y, but there is no need. \(y\ge 4x-2\) is in slope-intercept form, so there is no real need to do anything else to the equation. You can graph it pretty quick, or just try points. \(y\ge 4x-2\) for (0, −4) \((-4)\ge 4(0)-2\) \(-4\ge 0-2\) \(-4\ge -2\) Because -4 is less than -2, that is false and not an answer.

OpenStudy (accidentalaichan):

Oh, I see.

OpenStudy (accidentalaichan):

Thanks

OpenStudy (e.mccormick):

np. I'll also show the graphing method.

OpenStudy (e.mccormick):

To graph it, easy way is to find the intercepts: (0,?) and (?,0). Just use = for that. \(0= 4x-2\) \(2 = 4x\) \(2/4 = x\) \(1/2 = x\) So (1/2,0) is the x intercept \(y= 4(0)-2\) \(y= -2\) And (0,-2) is the y. I can draw a line through those two points. Because the question uses \(\ge\) it is a solid line.